Find the area under the curve y= x*3-6x*2 , the x-axis and the lines x =1 and x=2.
1
Expert's answer
2019-11-06T01:26:32-0500

Solution Using the properties of a definite integral, we obtain


"S=-\int_1^2(x^3-6x^2)dx=(2x^3-\frac {x^4} {4})|_1^2=16-4-2+\frac {1} {4}=\frac{41} {4}"

Answer

"\frac{41} {4}"